Understanding the Z-score



Solve the following problem.



1. X = 50, s = 5, X(mean) = 40


2. X = 40, o = 8, u(mean) = 52


3. X = 36, s = 6, X(mean) = 28

1
Expert's answer
2022-04-06T07:40:13-0400

1) "z=\frac{50-40}{5}=2."

2) "z=\frac{40-52}{8}=-1.5."

3) "z=\frac{36-28}{6}=1.33."